OpenStudy (pyro):

If you roll one dice and it lands on a 6, what is the percent chance of you rolling a 6 on another roll?

OpenStudy (anonymous):

It's always going to be the same probability, 1/6, because the events are independent and have nothing to do with what has transpired beforehand.

OpenStudy (pyro):

how would i turn that into a decimal?

OpenStudy (anonymous):

1/6 = 0.166666666...

OpenStudy (anonymous):

1/6=.16 .16*100=16%

OpenStudy (anonymous):

You can call that 0.166 with a bar over the last six which means it repeats.

OpenStudy (pyro):

thx u

OpenStudy (anonymous):

you're welcome!